Chapter 3 Mass Relationships in Chemical Reactions 6 Questions 12) What is the sum of the coefficlents when the following equation is balanced using the lowest whole-numbered coefficients? A) 10 13) Which 0) 36.5 Korc D) All of the A) 36.5 or c 14) What is the molar mass of aspartic acid, CaO4HN? C) 133 /mol D) 17 ol 0) 70 /mol A) 43 g/mof 1sI A compound responsible for the odor of garlic has a molecular weight of 7a efral 96e0 o.321 & of carbon, 0.044 of bydrogen, and .28ffur. What is the molecular formula A 0.650 e sample of the of the compound? 3 e/mol. C) C3H1553 D) C6H102 8) C3Hs5s 16) The following diagrams represent the reaction of A2 (shaded spheres) with B2 (unshaded spheres). Identify the limiting reactant and write a balanced equation for the reaction. A) CH5S 2 AB3. B) A2 is the limiting reactant; A2 + 3 B2 AJA2 is the limiting reactant; A + 3 B nt; A + 3 B AB3 AB3. 2 AB2 C) 82 is the limiting reactant, A+3B AB3 D)2 is the limiting reactant; A2+382 e? Molar masses:- CaCl2 111.0 g/mol a 17) How many grams of calcium chloride are needed to produce 10.0 g of potassium chlorid KCI 74.5 g/mol. CaCl2(aq) + K2CO3(aq)-) 2 Ka(aq) + CaCO3(aq) A) 0.134 g B) 7.44 g C) 9.28 g D) 18.56 g Chapter 4 Reactions in Aqueous Solutions 4 Questions 18) Write a balanced net ionic equation for the reaction of Pb(NO3)2laq) with Nalaq). B) Pb2+(aq)+21Haq)Pbl2(s) A) Pb(NO3)2(aq) + 2 Nal(aq) ? PbI2(s) + 2 NaNO3(aq) C) Pb2t(0q)+2 NO3-(aq) +2 Nat(oq) +2 I-(aq) Pbl2(s) + 2 Na+(oq)+2 NO3 (aq) D) Pb2t(oq) + 2 NO3 (aq)+ 2 Nat(aq)+2 1(aq) Pb2+(aq)+2(aq)+ 2 Nat(aq)+ 2 NO3 (ag) d o 16 g 19) What is the molar concentration of sodium ions in a 0.350 M NagPO? solution? (Molar masses; Na 23, P 31 a A) 0.117 M B) 0.350 M C) 1.05 MD 1.40 M 20) When dissolved in water, LiOH behaves as A) an acid that forms Lit and OH- ions. B) an acid that forms Lio- and H+ ions. C) a base that forms Lit and OH-ions. D) a base that forms Lio and Ht ions 21) The reaction 2 HNO3(aq)+ Ba(OH)2faq)Ba(NO3)2laq) + 2 H20(0 is best classified as a(n) A) acid-base neutralization reaction. C) precipitation reaction. B) oxidation-reduction reaction. D) single replacement reaction.
Answer #1

Ans 12 : C) 19

The reaction is balanced carefully noting down the elements on both the sides of the reaction . Then the elements on the left are equated with the elements on the right.

So the balanced reaction is given as :

4PH3 + 8O2 = P4O10 + 6H2O

So the sum of coefficients = 4 + 8 + 1 + 6 = 19